OpenStudy (anonymous):

Will someone please help me finish this problem: The solid is similar to a larger solid with the given scale factor, find the surface area and volume of the larger solid.cone: scale factor 3:4, SA=96pi cm^2, V=96pi cm^3

OpenStudy (anonymous):

so i have 96pi as the surface area and volume.. the new scale factor for sa is- 9/16 and for volume 27/64

OpenStudy (anonymous):

then what do i do to finish it

OpenStudy (anonymous):

Since you have your scale factors, your new Vol = <Old Volume> * <Volume Scaling factor>, and the same with S.A.

OpenStudy (anonymous):

so sa=96pi^2 and v=96pi^3

OpenStudy (anonymous):

SA = 96 pi^2 *(9/16)=54 pi^2 ; V = 96pi^2*(27/64) = (81/2) pi^2

OpenStudy (dumbcow):

SA = 96*(16/9) V= 96*(64/27)

OpenStudy (anonymous):

No, you cannot apply the scaling factor directly to area/volume because they don't change linearly with dimensions.